OpenStudy (anonymous):

how do i solve this? 8+10^(x)=1008

OpenStudy (shamim):

8+10^x=1008

OpenStudy (shamim):

10^x=1008-8=1000=10^3

OpenStudy (shamim):

now equate power

OpenStudy (anonymous):

where did the 3 come from?

OpenStudy (loser66):